A metal block having an internal cavity weighs 110 gms in air & 80 gms in water. If density of metal is 5.5gm/cm3 then find volume of cavity


 

Grade:12

1 Answers

Ramesh V
14 years ago

weight of object in water( = (Mass of object) *acceleration of gravity - Mass of an equal volume of water *acceleration of gravity.

density of water = 1gm/cc

Mw = Ma*g - Mv*g

80*10 = (110-Mv)*10

Mv = 30 gms

volume of cavity = 30cm3

--

regards

Ramesh